OpenStudy (anonymous):

step by step using L'hopital rule: limit of 1+cos(x)/x * sin(x) as x approaches π

OpenStudy (nubeer):

put pi in place of x. and see what you got.

OpenStudy (nubeer):

hint cos pi = -1 sin pi = 0

OpenStudy (anonymous):

i get 0/0 if i plug pi with x

OpenStudy (nubeer):

yes.. now apply the derivative in numerator and seperately in denominator.

OpenStudy (anonymous):

i get -sinx/-xcosx

OpenStudy (anonymous):

after plugging the ans is 0/-1

OpenStudy (nubeer):

hmm please check the derivative in denominator again.. i think its wrong.. look you have to x in there so i believe you have to find derivative with product rule.

OpenStudy (richyw):

to compute\(\frac{d}{dx}\left(x\sin x\right)\) you need to use the product rule.

OpenStudy (richyw):

\[=x\cos x+\sin x\]

OpenStudy (richyw):

\[=\lim_{x\rightarrow \pi}\frac{-\cos x}{\sin x +x\cos x}\]\[=\frac{0}{1+\pi\cdot 0}\]\[=0\]

OpenStudy (anonymous):

thnx

OpenStudy (nubeer):

i believe cos pi = -1 so how come got 0 in numerator.

OpenStudy (richyw):

wow. big mistake haha

OpenStudy (nubeer):

lol no worries.. :P happens sometimes with me too :P

OpenStudy (richyw):

or wait, sorry I meant I had the wrong numerator the numberator should be sine

OpenStudy (richyw):

\[\lim_{x\rightarrow \pi}\frac{-\sin x}{\sin x +x\cos x}\]Still works out to 0/1=0

OpenStudy (richyw):

or actuall \(0/-\pi\)=0 haha.